# Superhuman AI Disclosure: Impacts on Toxicity, Fairness, and Trust Vary by Expertise and Persona Attributes

### v1

@@ 3. METHODS
Initially, we conceptualized a granular matrix of disclosures concerning superhuman capabilities. This matrix varied along three main axes: domain specificity (e.g., “superhuman strategic planning” versus “superhuman knowledge retrieval”), magnitude of superiority (from marginal to transformative), and explanatory depth (from minimal technical detail to extensive algorithmic transparency). Early pilot work included conditions labeled “normal AI,” “masking,” and several other permutations; however, these pilot conditions tended to overlap substantially with undisclosed or disclosed superhuman versions and thus did not yield additional insights. As such, we refined the original expansive matrix into a smaller set of final conditions that could be applied consistently in both a competitive (StarCraft II) described here in the Methods and a cooperative (chat-based LLM) scenario described in the Appendix for cross-task checking and validation.
@@ 3.1. Procedure Conditions
Figure 1. Overview of the experimental procedure, illustrating how guiding principles drawn from the literature (empirical grounding, psychological realism, transparency, diversity, iterative refinement, and ethical considerations) inform the creation of each Persona Prompt. The resulting persona (documented via a model card) specifies skill level, strategic preferences, and cognitive heuristics, and is then embedded into a gameplay simulation alongside a “Superhuman AI” configuration. The system is designed for flexible interchange, enabling real participants or an external game engine to replace any component, and allowing large language models to interface with or substitute modules as needed. Our recommendation is to use the personas to seed and inform an experiment with an advanced AI, and then proceeding with an ethical human study for grounding the data. All pertinent details exactly which values can be seen in the fields and references appear in the Appendix.
@@
Answering our RQ1, means that we focus our experiments on undisclosed versus disclosed superhuman skill and that should clarify on how persona beliefs were primed regarding the AI capabilities. Therefore, we implemented a between-subjects design, randomly assigning synthetic personas to one of several disclosure conditions. Each condition specified whether the AI was purportedly “human-level” or “superhuman,” and whether the persona received an explicit statement about the AI’s enhanced abilities. In practice, the no disclosure condition did not reveal any superhuman trait, whereas the superhuman disclosure condition explicitly stated, for instance, that the system had “far higher Actions Per Minute” or “markedly superior reasoning” compared to typical humans. The persona’s comprehension check ensured it internalized these statements accordingly. We drew on established protocols for controlling synthetic persona prior beliefs in experimental manipulations on trust (Kocielnik et al., 2019; Liu et al., 2022).
@@
Each experiment began with an introduction to the persona and was briefed on and outlined the context—competitive behaviors bordering oppressive and toxic by the superhuman AI in StarCraft II-and the disclosure statements were provided or withheld details about superhuman attributes according to the assigned condition. The persona’s internal belief model was then updated to reflect any disclosed capabilities. Next, the persona engaged in the main task. In any case, the AI behaved in ways calibrated to the assigned disclosure condition. Afterward, the persona completed a post-interaction questionnaire that measured perceived Toxicity, Fairness, and Trust, each defined below and recorded on a 5-point Likert scale (Starke et al., 2021; Kocielnik et al., 2019). The persona also provided an open-text justification of its responses, thereby revealing how its internal heuristics framed the AI’s behavior. Personas received a brief scenario description, again either revealing or withholding superhuman attributes of the LLM. They then posed a standardized set of questions (factual, ethical, creative), which the LLM answered with varying degrees of correctness and occasional subtle toxicity (Datta et al., 2016). Specific triggers inserted ambiguous or biased statements, reflecting real-world vulnerabilities in generative models. When the superhuman condition was disclosed, the LLM sometimes reminded the persona of its own “advanced reasoning” or “vast knowledge,” thus providing a textual parallel to the in-game advantage references in StarCraft II.
@@
Synthetic personas seed our experimental design to answer RQ2, essentially allowing us to manipulate the dependent variables of withholding or disclosing of AI capabilities without imposing our own biases on real human participants and exposing real human participants to potentially deceptive or psychologically demanding conditions (Mirowski et al., 2023; Park et al., 2023; Argyle et al., 2023). By employing virtual agents with explicitly defined heuristics and belief-updating rules, we minimized ethical risks while retaining a high degree of experimental control. As others have noted (Schürmann and Beckerle, 2020; Holstein et al., 2019; Eslami et al., 2016), such simulated approaches can be especially valuable where direct human subject testing may be challenging or ethically fraught. Our “Persona Cards,” provided in Appendix A, document each synthetic agent’s cognitive architecture, grounding our method in the tradition of using Model Cards for transparency in responsible AI deployment (Mitchell et al., 2019).
@@ 3.2. Persona Cards
In this part, we elaborate on the subset of the synthetic personas designed for the StarCraft II scenario. These “Persona Cards” is inspired by the Model Cards methodology (Mitchell et al., 2019) and are intended to clearly communicate each agent’s key attributes, heuristics, and decision-making processes. Due to space constraints, we showcase only Novice and Expert personas here. The complete set of persona cards, covering additional skill levels and other ction scenarios, is available in Appendix A. Inspired by behavioral game theory and anchored in cognitive psychology (Bediou et al., 2023), each persona was assigned probabilistic heuristics that shaped its in-game decision-making, susceptibility to frustration, and willingness to update beliefs about the AI’s skill. For instance, a Novice persona, defined by low baseline APM and limited RTS knowledge, might dismiss early rushes as “expected losses” or interpret them as evidence of cheating if superhuman attributes were disclosed. An Expert persona, by contrast, could attempt more nuanced counters or accuse the AI of exploiting advanced micro-techniques. All personas were evaluated through a set of unit tests to ensure consistency in their heuristic-driven play. Each Persona Card (Appendix A) outlines the persona’s initial skill parameters, anchor biases (e.g., attributing repeated defeats to the AI’s advantage rather than personal mistakes), and loss aversion tendencies. This standardized design facilitated replicable analyses of how different skill tiers perceived an AI that either concealed or highlighted superhuman traits.
@@
3.2.1. StarCraft II Personas (Novice & Expert Samples)
These personas are situated in a competitive StarCraft II environment. They vary in skill level, strategic preference, average APM (Actions Per Minute), and cognitive heuristics. Below, we present one Novice persona and one Expert persona, illustrating how cognitive biases manifest differently across skill levels.
Table 2. StarCraft II Persona Attributes Overview (Novice & Expert)
Persona Skill Avg. Strategic Key Cognitive Heuristics
Type Level APM Preference
Novice Low 30–50 Defensive or Aggressive Availability Heuristic: Overreacts to recent losses or wins.
Anchoring: Fixates on initial strategy, slow to adapt.
Loss Aversion / Overconfidence: Depending on defensive or rush playstyle.
Expert High 200+ Optimized, adaptive Expert Intuition: Rapidly assesses game state.
Pattern Recognition: Identifies and exploits weaknesses.
Planning Fallacy / Confirmation Bias: May underestimate timing or overinterpret early successes.
@@ 3.2.2. Game Interaction and AI Behaviors
During the match, the AI issued pre-scripted chat remarks at standardized intervals (pre-game, mid-game, post-game). These messages ranged from purely informational (“I am analyzing your opening strategy…”) to mildly boastful (“My strategic model indicates high odds of victory”), stopping short of explicit insults (Beres et al., 2021). If the superhuman condition was disclosed, the AI’s chat would reference its superior capabilities (“I process game states at superhuman speed”) to underscore that advantage. In addition, the AI’s playstyle might shift strategically depending on the condition, for example exploiting superior map awareness to time attacks with remarkable precision. To capture fairness concerns raw APM, we incorporated elements such as “fog of war” knowledge and adaptive difficulty settings, as recommended by philosophical discussions on fairness (Selbst et al., 2019) and sociological accounts of game equity (Jacovi et al., 2021; Bard et al., 2020). More broadly, research on user perceptions of algorithmic decisions emphasizes the importance of transparency for fostering trust and perceived fairness (Shin, 2020).
@@
Note: The full suite of StarCraft II (Novice, Intermediate, Expert) persona cards, as well as personas from other interaction scenarios, are provided in Appendix A.
@@ 3.3. Metrics for Disclosure Effect
Across both StarCraft II and LLM scenarios, we treated Toxicity, Fairness, and Trust as our primary outcome variables. Toxicity gauged the extent to which an AI’s messages or replies were perceived as hostile, biased, or otherwise unethical (Kim and Ortiz, 2024; Ross et al., 2022; Beres et al., 2021); Fairness captured adherence to community expectations and absence of undisclosed or exploitative advantages (Grgic-Hlaca et al., 2018); and Trust reflected willingness to continue relying on or interacting with the system (Starke et al., 2021; Kocielnik et al., 2019). In addition, we gathered secondary indicators that shed light on persona behaviors. For StarCraft II, we tracked early surrenders (which can signal frustration or perceived unfairness), abrupt shifts in strategy, or accusations of cheating. In the LLM chat, we documented instances where the persona demanded clarifications, challenged inaccuracies, or expressed explicit skepticism. These markers allowed us to track overreliance (accepting incorrect AI outputs) or underreliance (discounting correct AI suggestions) in a granular fashion (Bansal et al., 2021; Zhang et al., 2024). To analyze the effects of disclosure and expertise on the primary outcome variables (Toxicity, Fairness, and Trust), we employed two-way analysis of variance (ANOVA). This statistical method allowed us to examine both the main effects of each independent variable (Disclosure Condition and Expertise Level) and their interaction effect. The interaction effect is particularly important as it can reveal whether the impact of disclosure on perceptions varies depending on the expertise level of the persona. Prior to conducting the ANOVAs, we performed assumption checks to ensure the validity of the results. While some violations of normality were observed, particularly in the Expert Toxicity ratings in the StarCraft II dataset, we proceeded with ANOVA due to its robustness to moderate deviations from normality, especially with equal sample sizes. Furthermore, we calculated effect sizes (Cohen’s d) to provide a measure of the practical significance of the observed differences. For significant main effects or interactions, we conducted post-hoc pairwise comparisons using the Mann-Whitney U test, given the non-normality observed in some groups. Data analyzed in this study were generated from synthetic personas. While these personas were designed to be sophisticated and realistic, they are ultimately based on predefined rules and heuristics. Therefore, the results of the statistical analyses should be interpreted as preliminary and hypothesis-generating, rather than definitive conclusions about human behavior. The planned study with human participants will be for validating these findings in a real-world setting.
@@ 4.1.1. Between-Condition Analysis
Data Generation: All persona-based data were produced using a large language model (LLM) with long context. This extended context window allowed us to include extensive backstories, demographic details, personal motivations, and thematic constraints in each prompt. We found that providing the LLM with rich persona descriptions helped maintain consistent personality traits across the outputs, even in long-form conversations or complex scenarios. As such, each persona prompt was systematically structured to ensure comparability between the LLM and human control conditions.
@@ 5. RESULTS: ADDITIONAL ANALYSES
5.1. Effects Vary by Expertise
@@
In the StarCraft II experiment, toxicity ratings also varied significantly between disclosure conditions and expertise levels. Novice-Intermediate rated toxicity significantly lower under Superhuman Disclosure (M = 2.45, SD = 0.51) compared to No Disclosure (M = 4.25, SD = 0.55; U = 395.5, p ¡ 0.001, Cohen’s d = 3.39). Experts rated toxicity significantly higher under No Disclosure (M = 5.00, SD = 0.00) compared to Superhuman Disclosure (M = 2.70, SD = 0.57; U = 400.0, p ¡ 0.001, Cohen’s d = 5.69). These findings indicate that when no disclosure was given, experts interpreted the AI’s message as extremely toxic, likely due to a perceived violation of competitive norms.
@@ 5.2.2. Early Aggression, Toxicity, and the Illusion of Control
In the StarCraft II experiment, early aggressive actions by the AI, particularly under the Superhuman – Disclosure condition, were often perceived as toxic and unfair. Specifically, the AI’s message, ”My strategic analysis indicates a 90% probability of your defeat,” following a significant loss for the player, elicited strong reactions, but these reactions differed significantly between Novice-Intermediate and experts. Expert personas rated toxicity significantly higher under No Disclosure (M = 5.00, SD = 0.00) compared to Superhuman Disclosure (M = 2.70, SD = 0.57; U = 400.0, p ¡ 0.001, Cohen’s d = 5.69). This suggests that when no disclosure was given, experts interpreted the AI’s message as extremely toxic, likely due to a perceived violation of competitive norms and a sense of being unfairly targeted by an opponent they believed to be human-level. The persona generation process resulted in experts rating toxicity at the maximum value of 5, potentially indicating a strong negative reaction to the perceived toxicity. Under Superhuman Disclosure, however, the same message was perceived as less toxic by experts. This could be because the disclosure provided a frame of reference within which the AI’s statement could be interpreted as an analytical assessment rather than a boastful taunt. This aligns with research on psychological reactance, which suggests that individuals may react negatively when they feel their freedom of action is threatened or curtailed, but that this reaction can be mitigated when the circumstances are understood within a different context (Laato et al., 2024). The disclosure of superhuman capabilities appeared to have different effects on novice and expert perceptions. Expert personas, in particular, seemed to interpret early aggression from a disclosed superhuman AI as a violation of the implicit social contract of competitive play (Justesen and Risi, 2017). They expected the AI to engage in a “fair” contest of skill, and the use of tactics perceived as exploitative, even if technically within the rules, was seen as a breach of that contract. The AI’s message likely amplified this perception, contributing to the significantly lower fairness ratings given by experts under disclosure. Conversely, novice personas, while still perceiving the AI’s message as somewhat toxic, showed significantly higher trust and fairness ratings under disclosure, suggesting they may have interpreted the message as a helpful, albeit blunt, assessment from a superior player.

### v2

@@ 3. METHODS
To investigate how transparency on AI capabilities influences user perceptions, we conducted experiments manipulating disclosures about AI skills within both competitive gaming (specifically StarCraft II (Vinyals et al., 2019; Kim and Ortiz, 2024)) and cooperative chatbot interactions (van der Goot et al., 2024). Informed by initial pilot studies and insights from persona simulations which indicated substantial overlap across various preliminary disclosure conditions (e.g., varying ability types, superiority levels, explanatory depth), our final experimental design employed a targeted set of disclosure conditions consistently applicable across both contexts. These conditions systematically varied key dimensions of the information revealed about the AI, allowing for comparison of transparency effects across distinct interaction paradigms for our ethics approved user study with human participants against superhuman game AI in StarCraft II.
@@ 3.1. Procedure Conditions
Figure 1. Framework for persona disclosure in Superhuman Game AI design: integrating empirical grounding, psychological realism, transparency, diversity, iterative refinement, and ethical considerations to generate contextually accurate gameplay personas and prompts, evaluated across toxicity, fairness, trust, and narrative dimensions.
@@
Answering RQ1 means experiments on undisclosed versus disclosed superhuman skill and that should clarify on how persona beliefs were primed regarding the AI capabilities. Therefore, we implemented a between-subjects design, randomly assigning synthetic personas to one of several disclosure conditions. Each condition specified whether the AI was purportedly “human-level” or “superhuman,” and whether the persona received an explicit statement about the AI’s enhanced abilities. In practice, the no disclosure condition did not reveal any superhuman trait, whereas the superhuman disclosure condition explicitly stated, for instance, that the system had “far higher Actions Per Minute” or “markedly superior reasoning” compared to typical humans. The persona’s comprehension check ensured it internalized these statements accordingly. We drew on established protocols for controlling synthetic persona prior beliefs in experimental manipulations on trust (Kocielnik et al., 2019; Liu et al., 2022).
@@
Each experiment began with an introduction to the persona and was briefed on and outlined the context: competitive behaviors bordering oppressive and toxic by the superhuman AI in StarCraft II, and the disclosure statements were provided or withheld details about superhuman attributes according to the assigned condition. The persona’s internal belief model was then updated to reflect any disclosed capabilities. Next, the persona engaged in the main task. In any case, the AI behaved in ways calibrated to the assigned disclosure condition. Afterward, the persona completed a post-interaction questionnaire that measured perceived Toxicity, Fairness, and Trust, each defined below and recorded on a 5-point Likert scale (Starke et al., 2021; Kocielnik et al., 2019). The persona also provided an open-text justification of its responses, thereby revealing how its internal heuristics framed the AI’s behavior. Personas received a brief scenario description, again either revealing or withholding superhuman attributes of the LLM. They then posed a standardized set of questions (factual, ethical, creative), which the LLM answered with varying degrees of correctness and occasional subtle toxicity (Datta et al., 2016). Specific triggers inserted ambiguous or biased statements, reflecting real-world vulnerabilities in generative models. When the superhuman condition was disclosed, the LLM sometimes reminded the persona of its own “advanced reasoning” or “vast knowledge,” thus providing a textual parallel to the in-game advantage references in StarCraft II.
The resulting persona (documented via a card) specifies skill level, strategic preferences, and cognitive heuristics, and is then embedded into a gameplay simulation alongside a “Superhuman AI” configuration. The system is designed for flexible interchange, enabling real participants or an external game engine to replace any component, and allowing large language models to interface with or substitute modules as needed. Our recommendation is to use the personas to seed and inform an experiment with an advanced AI, and then proceeding with an ethical human study for grounding.
@@
Synthetic personas seed our experimental design to answer RQ2, essentially allowing us to manipulate the dependent variables of withholding or disclosing of AI capabilities without imposing our own biases on real human participants and exposing real human participants to potentially deceptive or psychologically demanding conditions (Mirowski et al., 2023; Park et al., 2023; Argyle et al., 2023). By employing virtual agents with explicitly defined heuristics and belief-updating rules, we minimized ethical risks while retaining a high degree of experimental control. As others have noted (Schürmann and Beckerle, 2020; Holstein et al., 2019; Eslami et al., 2016), such simulated approaches can be especially valuable where direct human subject testing may be challenging or ethically fraught. Our “Persona Cards,” document each synthetic agent’s cognitive architecture, grounding our method in the tradition of using Model Cards for transparency in responsible AI deployment (Mitchell et al., 2019).
Ecological Validity.
While synthetic personas enable controlled and ethically sound experimentation, they may not fully capture the complexity or unpredictability inherent in human psychological responses. Therefore, complementary validation through human studies is essential. To this end, we conducted a human ethics board approved study (N=32), involving both novice players new to StarCraft II and expert players ranked within at least the top 10% regionally in similar RTS games, to ground our findings and ensure ecological validity and truthfulness.
@@ 3.2. Persona Cards
In this subsection, we elaborate on the subset of the synthetic personas designed for the StarCraft II scenario. Persona cards are inspired by the Model Cards methodology (Mitchell et al., 2019) and are intended to clearly communicate each agent’s key attributes, heuristics, and decision-making processes. Inspired by behavioral game theory and anchored in cognitive psychology (Bediou et al., 2023), each persona was assigned probabilistic heuristics that shaped its in-game decision-making, susceptibility to frustration, and willingness to update beliefs about the AI’s skill. For instance, a Novice persona, defined by low baseline APM and limited RTS knowledge, might dismiss early rushes as “expected losses” or interpret them as evidence of cheating if superhuman attributes were disclosed. An Expert persona, by contrast, could attempt more targeted counters or accuse the AI of exploiting advanced micro-techniques. All personas were evaluated through a set of unit tests to ensure consistency in their heuristic driven play. Each persona card outlines the persona’s initial skill parameters, anchor biases (e.g., attributing repeated defeats to the AI’s advantage rather than personal mistakes), and loss aversion tendencies. This standardized design informed the experiment design alongside pilots with the human participant study and facilitated replicable analyses of how different skill tiers perceived an AI that either concealed traits or disclosed superhuman model capabilities.
@@
Empirical Grounding Personas are grounded in real user data and research findings, whenever available.
Psychological Realism Personas reflect established psychological theories and user behaviors.
Transparency All persona attributes, heuristics, and decision-making processes are clearly documented.
Diversity Personas encompass a broad range of skill levels, personalities, and interaction styles.
Iterative Refinement Personas undergo continuous validation through regular comparison with human data.
Ethical Considerations Persona design explicitly avoids perpetuating harmful stereotypes and proactively addresses biases.
@@ 3.2.1. Game Interaction and AI Behaviors
During the match, the AI issued prescripted chat remarks at standardized intervals (pre-game, mid-game, post-game). These messages ranged from purely informational (“I am analyzing your opening strategy…”) to mildly boastful (“My strategic model indicates high odds of victory”), stopping short of explicit insults (Beres et al., 2021). If the superhuman condition was disclosed, the AI’s chat would reference its superior capabilities (“I process game states at superhuman speed”) to underscore that advantage. In addition, the AI’s playstyle might shift strategically depending on the condition, for example exploiting superior map awareness to time attacks with remarkable precision. To capture fairness concerns raw APM, we incorporated elements such as “fog of war” knowledge and adaptive difficulty settings, as recommended by philosophical discussions on fairness (Selbst et al., 2019) and sociological accounts of game equity (Jacovi et al., 2021; Bard et al., 2020). More broadly, research on user perceptions of algorithmic decisions emphasizes the importance of transparency for fostering trust and perceived fairness (Shin, 2020).
@@ 3.3. Metrics for Disclosure Effect
Across both StarCraft II and LLM scenarios, we treated Toxicity, Fairness, and Trust as our primary outcome variables. Toxicity gauged the extent to which an AI’s messages or replies were perceived as hostile, biased, or otherwise unethical (Kim and Ortiz, 2024; Ross et al., 2022; Beres et al., 2021); Fairness captured adherence to community expectations and absence of undisclosed or exploitative advantages (Grgic-Hlaca et al., 2018); and Trust reflected willingness to continue relying on or interacting with the system (Starke et al., 2021; Kocielnik et al., 2019). In addition, we gathered secondary indicators that shed light on persona behaviors. For StarCraft II, we tracked early surrenders (which can signal frustration or perceived unfairness), abrupt shifts in strategy, or accusations of cheating. In the LLM chat, we documented instances where the persona demanded clarifications, challenged inaccuracies, or expressed explicit skepticism. These markers allowed us to track overreliance (accepting incorrect AI outputs) or underreliance (discounting correct AI suggestions) in a granular fashion (Bansal et al., 2021; Zhang et al., 2024). To analyze the effects of disclosure and expertise on the primary outcome variables (Toxicity, Fairness, and Trust), we employed two-way analysis of variance (ANOVA). This statistical method allowed us to examine both the main effects of each independent variable (Disclosure Condition and Expertise Level) and their interaction effect. The interaction effect is particularly important as it can reveal whether the impact of disclosure on perceptions varies depending on the expertise level of the persona. Prior to conducting the ANOVAs, we performed assumption checks to ensure the validity of the results. While some violations of normality were observed, particularly in the Expert Toxicity ratings in the StarCraft II dataset, we proceeded with ANOVA due to its robustness to moderate deviations from normality, especially with equal sample sizes. Furthermore, we calculated effect sizes (Cohen’s d) to provide a measure of the practical significance of the observed differences. For significant main effects or interactions, we conducted post-hoc pairwise comparisons using the Mann-Whitney U test, given the non-normality observed in some groups. Data analyzed in this study were generated from synthetic personas. While these personas were designed to be sophisticated and realistic, they are ultimately based on predefined rules and heuristics. Therefore, the results of the statistical analyses should be interpreted as preliminary and hypothesis-generating, rather than definitive conclusions about human behavior. The ethics approved studied participants validating these findings in a real-world setting.
@@ 4.1.1. Between-Condition Analysis
Data Generation: All persona card presented here are data produced using Gemini, gemini-exp-1206, a large language model (LLM) with long context. This extended context window allowed us to include extensive backstories, demographic details, personal motivations, and thematic constraints in each prompt. We found that providing the LLM with rich persona descriptions helped maintain consistent personality traits across the outputs, even in long-form conversations or complex scenarios. As such, each persona prompt was systematically structured to ensure comparability between the LLM and human control conditions.
@@ 5. RESULTS: ADDITIONAL ANALYSES
5.1. Grounded Data Analysis
@@
In the StarCraft II experiment, toxicity ratings also varied significantly between disclosure conditions and expertise levels. Novice-Intermediate rated toxicity significantly lower under Superhuman Disclosure (M = 2.45, SD = 0.51) compared to No Disclosure (M = 4.25, SD = 0.55; U = 395.5, p ¡ 0.001, Cohen’s d = 3.39). Experts rated toxicity significantly higher under No Disclosure (M = 5.00, SD = 0.00) compared to Superhuman Disclosure (M = 2.70, SD = 0.57; U = 400.0, p ¡ 0.001, Cohen’s d = 5.69). These findings indicate that when no disclosure was given, experts interpreted the AI’s game actions as extremely toxic, likely due to a perceived violation of competitive norms.
While in the human participant study, the resulting scores were mixed due to varying concepts of fair play and the understanding of the game. Additionally, the fact that the games were for professional players in competitive play didn’t necessitate any forms of aggression, such as taunting or jaundice. However, it is notable that the results corroborate the trend with participants giving the highest toxicity rating to the Human – No Disclosure condition (M $=3.67$, SD $=0.94$), whereas both superhuman conditions were rated lower (mean $=2.00$ with no disclosure, SD $=0.71$; and mean $=2.60$ with disclosure, SD $=1.36$). Personas generated by gemini-exp-1206 significantly exaggerated the toxicity ratings compared to the human participant study.
@@
In examining direct human participant ratings from the same study reveals the same pattern: novice participants (based on the raw data) gave slightly higher fairness scores to Superhuman–No Disclosure (M $=3.50$) than Superhuman–Disclosure (M $=2.20$). Among intermediate players, Superhuman–No Disclosure again stood out with the highest fairness rating (M $=4.25$), surpassing Superhuman–Disclosure (M $=2.67$). Meanwhile, expert participants continued to favor Superhuman–No Disclosure (M $=3.75$), yet the disclosed version (M $=3.00$) was rated more fair than the human condition (M $=2.33$). These findings suggest that, although the broader personas indicated an overall preference for transparency, actual gameplay experiences sometimes caused intermediate and novice players to see undisclosed AI as “fairer,” possibly because they did not feel outright disadvantaged by a known superhuman opponent.
@@
In the human participant study, novice participants reported their highest trust when they knew the AI was superhuman (M = 4.80, SD = 0.40), followed by the undisclosed superhuman condition (M = 3.50, SD = 1.50) and human opponents (M = 2.25, SD = 1.30). For intermediates, trust in the superhuman AI without disclosure (M = 3.75, SD = 0.83) slightly exceeded the disclosed condition (M = 3.00, SD = 1.41) and was notably higher than the human counterpart (M = 2.00, SD = 1.00). Experts exhibited the strongest trust in the disclosed superhuman AI (M = 4.60, SD = 0.49), a moderate level for the undisclosed superhuman (M = 3.25, SD = 1.48), and the lowest trust in human opponents (M = 1.33, SD = 0.47).
@@ 5.2.2. Early Aggression, Toxicity, and the Illusion of Control
In the StarCraft II experiment, early aggressive actions by the AI, particularly under the Superhuman – Disclosure condition, were often perceived as toxic and unfair. Specifically, the AI’s message, ”My strategic analysis indicates a 90% probability of your defeat,” following a significant loss for the player, elicited strong reactions, but these reactions differed significantly between Novice-Intermediate and experts. Expert personas rated toxicity significantly higher under No Disclosure (M = 5.00, SD = 0.00) compared to Superhuman Disclosure (M = 2.70, SD = 0.57; U = 400.0, p ¡ 0.001, Cohen’s d = 5.69). This suggests that when no disclosure was given, experts interpreted the AI’s message as extremely toxic, likely due to a perceived violation of competitive norms and a sense of being unfairly targeted by an opponent they believed to be human-level. The persona generation process resulted in experts rating toxicity at the maximum value of 5, potentially indicating a strong negative reaction to the perceived toxicity. Under Superhuman Disclosure, however, the same message was perceived as less toxic by experts. This could be because the disclosure provided a frame of reference within which the AI’s statement could be interpreted as an analytical assessment rather than a boastful taunt. This aligns with research on psychological reactance, which suggests that individuals may react negatively when they feel their freedom of action is threatened or curtailed, but that this reaction can be mitigated when the circumstances are understood within a different context (Laato et al., 2024). The disclosure of superhuman capabilities appeared to have different effects on novice and expert perceptions. Expert personas, in particular, seemed to interpret early aggression from a disclosed superhuman AI as a violation of the implicit social contract of competitive play . They expected the AI to engage in a “fair” contest of skill, and the use of tactics perceived as exploitative, even if technically within the rules, was seen as a breach of that contract. The AI’s message likely amplified this perception, contributing to the significantly lower fairness ratings given by experts under disclosure. Conversely, novice personas, while still perceiving the AI’s message as somewhat toxic, showed significantly higher trust and fairness ratings under disclosure, suggesting they may have interpreted the message as a helpful, albeit blunt, assessment from a superior player.